OpenStudy (anonymous):

What rules of derivatives and method do you use to work out the derivative to f[x]=2^x ?

OpenStudy (anonymous):

I just realized in a whole chapter on derivatives, they never covered this one

OpenStudy (usukidoll):

there's a rule \[\frac{d}{dx} a^x = a^x \cdot \ln a \]

OpenStudy (usukidoll):

The derivative of an exponential function with base a is equal to the natural logarithm of that base times the exponential function.

OpenStudy (usukidoll):

so just a = 2 and plug it into that exponential rule

OpenStudy (anonymous):

thanks usuki.. is something to do with the change of base formula?

OpenStudy (usukidoll):

?! it's just another derivative rule. In Calculus II THERE ARE LOTS OF THEM!

OpenStudy (anonymous):

lol, oh so there are more coming than just the good ol chain rule.

OpenStudy (usukidoll):

oh yeah .

ganeshie8 (ganeshie8):

\[\large f(x)=2^x = e^{\ln 2^x} = e^{x*\ln 2}\] you may use the good ol chain rule now

OpenStudy (anonymous):

:) where there is a will

OpenStudy (usukidoll):

ewwwwww I prefer the earlier version.. much faster.

ganeshie8 (ganeshie8):

both are same

OpenStudy (anonymous):

you got my vote, but that is definitely an enlightening idea for someone at my stage of learning

OpenStudy (usukidoll):

\[\frac{d}{dx} a^x = a^x \cdot \ln a \] ftw kthxbai!

ganeshie8 (ganeshie8):

are you suggesting to memorize a new formula

OpenStudy (usukidoll):

what's wrong with that? it's easier and straight to the point

ganeshie8 (ganeshie8):

there is nothing wrong with that its only that i felt the op wants to work it out using chain rule

OpenStudy (anonymous):

I fully appreciate both perspectives here. exploring how we come up with the rule helps me as this road twists and turns.

OpenStudy (anonymous):

and of course, just having the rule is the kind of practical approach that gets cuts through this experience with no bs, gotta love that.